# Meta Pruning via Graph Metanetworks : A Meta Learning Framework for Network Pruning

### v1

@@ 3 Method
3.1 Equivalent Conversion between Neural Network and Graph
To enable a neural network to be processed by another network, we first need to represent the original network in a structured data format. In this work, we convert the neural network into a graph, drawing inspiration from prior works such as Kofinas et al. [30], Lim et al. [40], but with notable differences in implementation.
To convert a network into a graph, we establish the correspondence between the components of a neural network and elements of a graph as follows:
  • Node: Each neuron or convolutional channel in the network is represented as a node in the graph.
  • Edge: An edge exists between two nodes if there is a direct connection between them in the original network. This includes connections between neurons or channels within linear or convolutional layers, as well as residual connections.
  • Node Features: Node features comprise parameters associated with neurons, here we use the weight, bias, running mean, running var of batchnorm [26], including both batchnorms from previous adjacent layers and previous skip connection layers (if exist).
  • Edge Features: Edge features encode the weights of connections between nodes. We treat linear connections and residual connections as special cases of convolutional connections with a kernel size of 1. For a $k\times k$ convolution between two channels, we linearize it into a $k^{2}$-dim edge feature vector.
Since graph neural networks require all edge features to have the same dimension, we adopt one of two strategies to standardize them:
  • Pad all convolutional kernels to the same size (e.g., the maximum kernel size in the network), then flatten them to form fixed-length edge feature vectors.
  • Flatten convolutional kernels of varying sizes, then apply learned linear transformations to project them into a uniform feature dimension.
Figure 1: We visualize a two layer small CNN as example. Our graph contains 6 nodes corresponding with 6 neurons in the network. For neurons connected via convolutional channel like (1,3) and connected by residual connection like (1,5), we add an edge between their corresponding nodes. Then we generate node and edge features as shown in the figure. For some special cases, we use default values to replace non-existent values. For example, for node 3 who doesn’t have a residual connection, we set the last 4 number of its node feature as [1, 0, 0, 1]. This means we treat it as with previous residual batchnorm of weight 1, bias 0, running mean 0, running var 1, performing the same as with no batchnorm.
3.2 Graph metanetwork design
@@
After changing the input network into a graph, we use our metanetwork, a graph neural network, to process the input graph. We build our metanetwork based on the message passing framework [12] and PNA [3] architecture.
@@
We multiply our predictions by a residual coefficient and add them to the original inputs as our outputs:
@@
3.3 Group $l_{2}$ Norm Criterion
We use group $l_{2}$ norm from Fang et al. [9] as our pruning criterion. The design philosophy of group $l_{2}$ norm is to measure the sparsity of pruning groups. In structural pruning, we only prune a group of elements in a network. For example, if we prune a neuron, then all edges connected to the neuron should also be pruned, which we say that they are in a group. A more detailed explanation of group can be found in Fang et al. [9]. Since structural pruning is performed on groups, traditional $l_{2}$ norm that treats nodes and edges in isolation is no longer viable. Group $l_{2}$ norm that takes a whole group into consideration has shown to be more effective.
Given a parameter group $g=\{w_{1},w_{2},\ldots,w_{|g|}\}$. For each parameters $w$ with $K$ prunable dimensions indexed by $w[k]$, we measure the importance of the k-th prunable dimension by :
$I_{g,k}=\sum_{w\in g}\|w[k]\|^{2}$ (13)
@@ 3.3 Group $l_{2}$ Norm Criterion
$\mathcal{R}(g,k)=\sum_{k=1}^{K}\gamma_{k}\cdot I_{g,k}=\sum_{k=1}^{K}\sum_{w\in g}\gamma_{k}\|w[k]\|^{2}$ (14)
@@
$\gamma_{k}=2^{\alpha(I_{g}^{\max}-I_{g,k})(I_{g}^{\max}-I_{g}^{\min})}$ (15)
And we use $\alpha=4$ in all our experiments.
When pruning, we calculate the normalized relative score of all prunable dimensions in all groups
$\hat{I}_{g,k}=N\cdot I_{g,k}/\sum\{\text{TopN}(I_{g})\}$ (16)
where $N$ is a hyperparameter that chooses the top $N$ importance scores in the group $g$ to normalize every prunable dimension $k$’s score $I_{g,k}$. We prune unimportant dimensions of groups globally from lower to higher according to the normalized relative score.
Figure 2: Meta-Training Pipeline.
3.4 Meta-Training Pipeline
The goal of training our metanetwork to enable it to transform a hard-to-prune network into an easy-to-prune one, referred to as meta-training. This section details our meta-training pipeline, which is illustrated in Figure 2.
To prepare for meta-training, first we need the training data. Meta-training data is comprised of two parts. The first is a set of neural networks to prune, typically between 2 to 8, which serve as the origin network in meta-training. The second is the traditional training datasets, such as CIFAR-10, CIFAR-100, or ImageNet, used to calculate the accuracy loss.
During each training iteration, we select one model from data models as the origin network. The origin network is then converted into its graph representation. This graph is passed through the metanetwork, which outputs a new graph. We subsequently convert this new graph back into a new network.
Two types of losses are computed on the new network:
1) Accuracy Loss: We feed the training data into the new network and compute the cross-entropy loss based on the output predictions, ensuring the new network to have high accuracy after finetuning.
2) Sparsity Loss: We apply the group $l_{2}$ norm (as defined in 14) on the new network to encourage structured sparsity, making it more amenable to pruning.
The gradients from the loss are backpropagated through the metanetwork to update its parameters.
The accuracy loss ensures that the metanetwork preserves the performance of the original network, while the sparsity loss encourages the metanetwork to generate a network with group sparsity that facilitates pruning. After meta-training, our metanetwork can generate a transformed network that has almost no loss in accuracy after finetuning, and is significantly easier to prune.
Empirically, we observe that the metanetwork performs much better on hard-to-prune origin networks. However, most networks are initially easy to prune—for example, ResNet-56 on CIFAR-10 can be pruned with a 1.3× speedup and finetuned without accuracy loss. Therefore, in all our experiments (both during meta-training and pruning), we first prune the network slightly—without any accuracy degradation—to make it harder to prune. This pre-pruning step allows us to fully leverage the capability of the metanetwork.
3.5 Pruning a Target Network
Our pruning pipeline only involves a feedforward through metanetwork and finetuning. It can be summarized as:
@@ 3.5 Pruning a Target Network
$InitialPruning(Finetuning)\rightarrow Metanetwork(Finetuning)\rightarrow Pruning(Finetuning)$
As discussed in 3.4, we first do initial pruning to make our network harder to prune while maintaining the accuracy. Then we feed the hard-to-prune network through the metanetwork to transform it into a much easier-to-prune one, which also has almost no accuracy loss after finetuning. In the end, we perform the final pruning to reduce its size, and the final finetuning to evaluate the performance.
4 Experiments
4.1 Setup
We carry out our experiments on three most classical, popular, and representative image recognition tasks, including resent56 [16] on CIFAR10 [31], VGG19 [54] on CIFAR100 [31] and resnet50 [16] on ImageNet [4]. See appendix A for more general setups.

### v2

@@ 6.2 Transfer between Datasets and Architectures
CIFAR10 93.35 92.47 92.87 91.28
CIFAR100 69.97 70.16 69.25 68.91
SVHN 96.79 96.50 96.86 96.78
Transfer between architectures: 2 architectures, ResNet56 and ResNet110 (He et al., 2016) are used. We train metanetwork and the to be pruned network on each of them, traverse through every possible combination, and use the metanetwork to prune the network. Results (Table 4) show that our metanetwork can transfer between similar datasets. See Appendix F.2 for full details
Table 4: Transfer between architectures. All training dataset is CIFAR10. Columns represent the architectures used for training the metanetwork, and rows represent the architecures of the to be pruned network. “None” indicates using no metanetwork. All results with metanetwork is obviously better than no metanetwork.
[…]
ResNet56 93.40 92.81 92.08
ResNet110 93.04 93.38 92.40
Possible Future Use: For large-scale networks, the metanetwork can be pretrained on smaller architectures of similar design. And when the original training dataset is unavailable or excessively large, a related dataset may be used for pretraining. This transferability substantially enhances the practicality of the framework.
7 Experiments on Transformers
As mentioned in section 4.4, our framework is theoretically applicable to almost all types of networks with all kinds of pruning. Here we expand our implementations to another widely used arthitecture–transformer (Vaswani et al., 2017). More specifically, we pruned the vision transformer (ViT) from Dosovitskiy et al. (2021) that is trained on ImageNet. See Appendix G for full experiments, including how we convert transformer into graph and conduct further meta-training and pruning.
8 Conclusion
We propose an entirely new meta-learning framework for network pruning. For a pruning criterion, we use a metanetwork to change a hard to prune network into another easy to prune network for better pruning. This is a general framework that can be theoretically applied to almost all types of networks with all kinds of pruning. We present practical implementations of our framework and achieve outstanding results. Further analysis and experiments show that our framework has natural generality, flexibility, and transferability.
9 Reproducibility statement
Our code is available at https://anonymous.4open.science/r/MetaPruning together with all the guides to reproduce our experiments.
Appendix A Comparison with prior works
Our framework is entirely new and fundamentally different from all prior works. In this section, we will compare our method with earlier ones to help you better understand the relationship between our method and prior ones, and realize our unique innovations and advantages.
A.1 General Comparison
Previous pruning approaches can be broadly categorized into fixed pruning methods and learning to prune methods. Our approach falls into the learning to prune category but learns in a completely different way. For clarity, we provide a detailed comparison in Table 5 to help you better understand how our work relates to previous studies.
Table 5: Our framework is entirely new and has many advantages over previous works
Fixed pruning Learning to prune Ours
Definition Pruning with fixed hand-crafted algorithms. Using neural network learning techniques to learn how to prune. Also learning to prune, but learning in a completely different way.
Performance Bad. Networks are complex and hand-crafted algorithms are quite limited Good. Good (almost best according to our experiments).
Cost Low. No special extra training needed. High. Need special extra training during each pruning. Need special extra training before pruning. But once the training is done, can prune as many networks as we want without any special extra training.
Potential Has already been well-studied and commonly used. The pruning process is too tricky and costly. Can only be used in a very specific situation and lack of generality. The pruning process is relatively more efficient. Can theoretically be used in almost all situations and has great generality.
A.2 Computational and Memory Costs
One important question for almost all previous meta-learning approaches is the computational and memory costs are too large. We provide some estimate of memory and computational costs for scaling our methods to larger models and compare with two classic pruning methods–pruning at initalization methods (cheap pruning, hard to training and low accuracy like Lee et al. (2019); Wang et al. (2022)) and iterative magnitude pruning (costly, pruning during training and high performance like Frankle & Carbin (2019); Molchanov et al. (2019)). The results show that our method achieves outstanding results while uses relatively low costs.
For convenience in expression, we refer to pruning at initialization as “init pruning”, iterative magnitude pruning as “iter pruning” and our method as “meta pruning”.
For all usual networks which satisfy Edge Number >> Neuron Number , the memory and computational cost are all $O(E)$ ($E=$ Edge Number). This is obvious for init and iter pruning. For meta pruning, our metanetwork doesn’t scale with the network. So all computation and memory cost is it still $O(E)$.
PS: The reason why our metanetwork doesn’t scale with the network is that it only learns how to change weights based on local architectures and weights (the range of “local” is fixed for network of different sizes). In all our experiments, it is a small network compared to the to be pruned network, and it works well. So we have the confidence that it shouldn’t and doesn’t need to scale with the networks.
While all methods share the same asymptotic cost, their constant factors differ significantly.
Computational cost breaks down as:
$\text{Compu Cost}=\text{Training cost}+\text{Pruning cost}+\text{Other cost}$
  • Init pruning: Low training and pruning costs, negligible other cost.
  • Iter pruning: High training cost (initial training + multiple finetuning steps) and high pruning cost (multiple pruning rounds), no extra cost.
  • Meta pruning: Moderate training cost (initial training + 3 finetuning steps), moderate pruning cost (2 pruning steps), plus a small other cost consisting of metanetwork feedforward and amortized meta-training cost per network.
We estimate the relative magnitude in the following table:
Computation Training cost Pruning cost Other cost
Init 1T 1P 0
Iter >10>10T >5>5P 0
[…]
Where:
T: Unit training cost
P: Unit pruning cost
[…]
N: Number of networks pruned; meta-training cost is amortized over N (Once we get a metanetwork by meta-training, we can use it to prune as many networks as we want)
A: Meta-training cost. A is estimated 5–20.
Memory cost depends on both the amount of memory used and the duration of usage:
Method Amount Duration All
Init Small Short Low
Iter Medium Long Large
Meta Large for a very short time, Medium Medium but requires
Small rest of the time high memory capacity
  • Init pruning: Low memory usage for a short time.
  • Iter pruning: Medium memory usage for a very long time.
  • Meta pruning: The “feed forward through metanetwork” process requires large memory usage but takes a very short time. For the rest of the time it uses little memory. It is faster than iter pruning but slower than init pruning. While the average memory cost is moderate, peak usage demands high memory capacity.
Meta-training requires large memory usage for a long time. But its cost can be amortized into each pruning like we mentioned before. Take pruning resnet50 or ViT on ImageNet as example (which is already quite large model and dataset), NVIDIA A100 with 80GiB VRAM is enough for meta-training and feedforward through metanetwork and the rest training and finetuning can be done on NVIDIA RTX 4090 with 24 GiB VRAM. When we don’t have enough memory capacity, we can also change the batch size and use more time to make up for the lack of our memory capacity.
In summary, our method achieves outstanding results while uses relatively low costs.
A.3 Generality
Almost all previous meta-learning based pruning approaches, such as Liu et al. (2019); Li et al. (2020); Wu et al. (2024), are tailored to a specific network and thus lack generality. Because of this they require specialized training for each pruning instance.
Unlike prior methods, our metanetwork learns a universal rule for adapting weights based on local information, thereby enabling more effective pruning without relying on layer-specific or architecture-specific heuristics. We need no special training during each pruning. Once our metanetwork is trained, it can be used to prune as many networks as we want, and even transfer between datasets and architectures.
Appendix B Framework implementation details
B.1 Metanetwork(GNN) Architecture
@@
We build our metanetwork based on the message passing framework (Gilmer et al., 2017) and PNA (Corso et al., 2020) architecture.
Notation:
$a\leftarrow b$ : $\quad\text{: assignment/update, overwrite }a\text{ with the value of }b$ (10)
$X\odot Y$ : $\quad\text{: Hadamard/elementwise product, }(X\odot Y){i}=X_{i}Y_{i}$ (11)
@@
We multiply predictions by a residual coefficient and add them to the original inputs as our final outputs:
@@
B.2 Pruning Criterion
B.2.1 DepGraph and Torch-Pruning
DepGraph (Fang et al., 2023) proposes a general sparsity regularization based structural pruning framework that can be applied to a wide range of neural network architectures, including CNNs, RNNs, GNNs, Transformers, etc. Alongside the paper, the authors released Torch-Pruning, a powerful Python library that enables efficient structural pruning for most modern architectures. In our work, the pruning criterion is designed based on the methodology of DepGraph (Fang et al., 2023), and all pruning operations are implemented using Torch-Pruning.
Figure 7: A picture from (Fang et al., 2023). It shows how we group parameters together in structural pruning.
B.2.2 Structural pruning
In all our experiments, we only do structural pruning. Although we did not conduct experiments with unstructured pruning, our implementations are directly applicable to that setting as well. Structural pruning involves removing parameters in predefined groups, so that the resulting pruned model can be used as a standalone network—without relying on the original model with masks, nor requiring specialized AI accelerators or software to realize reductions in memory footprint and computational cost.
In structural pruning, a pruning group consists of all parameters that must be pruned together to maintain network consistency. For example, in Figure 7(a), if an input channel of convolution layer $f_{4}$ is pruned, the corresponding channel in batch normalization (BN) layer $f_{2}$ and the output channel of convolution layer $f_{1}$ must be pruned as well. Furthermore, due to the presence of a residual connection, the corresponding channels in BN layer $f_{5}$ and the output channel of convolution layer $f_{4}$ must also be removed. All of these channels together form a single pruning group. We define the number of prunable dimensions of a group as the size of the input channel dimension of Conv $f_{4}$ (which is equivalent to the corresponding dimensions in BN $f_{2}$, the output channel of Conv $f_{1}$, etc.)
For a more comprehensive theory to find all pruning groups in structural pruning, refer to the DepGraph (Fang et al., 2023) paper (section 3.1 & 3.2).
B.2.3 A importance score function
Given a parameter group $g=\{w_{1},w_{2},\ldots,w_{|g|}\}$ with K prunable dimensions indexed by $w_{t}[k]$ ($t\in\{1,2...|g|\}$), the score of the k th prunable dimension in group g is written is $I_{g,k}$, we introduce a general way to generate a series types of importance scores. The way has two key concepts, REDUCE and NORMALIZE. We first use REDUCE to reduce scores of parameters in the same group and in the same prunable dimension into one score.
$I_{t,k}=\mathop{\text{REDUCE}}\limits_{t:t\in\{1,2...|g|\}}(|w_{t}[k]|^{p})$ (24)
Here p is a hyperparameter and we usually set it to 2, REDUCE can be (but is not limited to) the following:
  1. (1)
    MEAN: $(w_{1}+w_{2}+...+w_{|g|})/|g|$
  2. (2)
    FIRST: $w_{1}$
Then we use NORMALIZE to normalize scores in the different prunable dimensions of the same pruning group.
$\hat{I}_{g,k}=I_{g,k}/\mathop{\text{NORMALIZE}}\limits_{k:k\in\{1,2...K\}}(I_{g,k})$ (25)
Here NORMALIZE can be (but is not limited to) the following:
  1. (1)
    NONE: 1 (use no normalize).
  2. (2)
    MEAN: $(I_{g,1}+I_{g,2}+...+I_{g,K})/K$
  3. (3)
    MAX: The maximum amoung $I_{g,1},I_{g,2}...I_{g,K}$
We calculate scores of all prunable dimensions in all groups in a network, rank them and prune according to their scores from lower to higher.
B.2.4 A sparsity loss
@@
$\mathcal{R}(g,k)=\sum_{k=1}^{K}\gamma_{k}\cdot\hat{I}_{g,k}$ (26)
@@
$\gamma_{k}=2^{\alpha\frac{\sqrt{I_{g}^{\max}}-\sqrt{I_{g,k}}}{\sqrt{I_{g}^{\max}}-\sqrt{I_{g}^{\min}}}}$ (27)
Here $\alpha$ can be but is not limited to:
  1. (1)
    4: What we use as default in all our experiments.
  2. (2)
    0: Same as using no shrinkage strength.
In sparsity loss, we treat $\gamma_{k}$ and NORMALIZE in the denominator of $\hat{I}_{g,k}$ simply as constants, which means gradients aren’t backpropagated through them. Gradients only back propagated from $\mathcal{R}(g,k)$ to $I_{g,k}$ to the parameters of the network.
In our code, rather than calculate the sparsity loss, we directly modify the gradients of the parameters of the network, which get the same results.
B.2.5 Kinds of pruning criterion
By choosing different ways of REDUCE and NORMALIZE (appendix B.2.3) and different $\alpha$ (appendix B.2.4), we can make different pruning cirteria in a uniform framework. In most of our experiments, we use MEAN REDUCE, MAX NORMALIZE, $\alpha=4$, as our default pruning criterion, and we name it group $\ell_{2}$ norm max normalizer. We also tried different criteria in our experiments and found they are almost all effective, which further shows the flexibility and effectiveness of our framework.
B.2.6 Pay Attention
Though we draw ideas from Fang et al. (2023) and use their Torch-Pruning libirary, we find their paper is inconsistent with their code. We wrote appendix B.2.3 & B.2.4 strictly based on their code, which is the Torch-Pruning library. If you read the Fang et al. (2023) paper section 3.3, you may find our description a little different from theirs, because their paper is inconsistent with their code, and we are in line with their code.
Appendix C Experimental Preliminaries
C.1 Terminologies
Speed Up: In line with prior works, we define $\textbf{Speed Up}\triangleq\textbf{Origin FLOPs }/\textbf{ Pruned FLOPs}$. It reflects the extent to which our pruning reduces computation, thereby accelerating the network’s operation.
Acc vs. Speed Up Curve: We generate the curve by prune the network little by little. After each pruning step, we evaluated the model’s performance on the test set (no finetuning in this step) and recorded a data point representing the current speed up and corresponding test accuracy. Connecting these points forms a curve that illustrates how accuracy changes as the speed up increases.
C.2 Pruning Pipeline
Our pruning pipeline can be summarized as:
@@ C.2 Pruning Pipeline
$\underbrace{\text{Initial Pruning (Finetuning)}}_{\text{Optional}}\rightarrow\underbrace{\text{Metanetwork (Finetuning)}\rightarrow\text{Pruning (Finetuning)}}_{\text{Necessary}}$ (28)
Initial Pruning: Empirically, we observe that the metanetwork performs better on hard to prune origin networks. However, many networks are initially easy to prune. For instance, ResNet56 on CIFAR10 can be pruned with a $1.3\times$ speedup and then finetuned without any accuracy loss. Therefore, unless stated otherwise, we apply the initial pruning step in all our experiments—both when generating meta-training data models and during pruning. In this step, we slightly prune the original network with a fixed speed up (a hyperparameter that can be determined in a few quick trials) to make it harder to prune without loss in accuracy (Figure 8(a)), thereby exploiting the full potential of the metanetwork. This effect is analogous to removing low-quality samples from a dataset in order to obtain a higher-quality subset: an easy to prune network contains many low-quality parameters, which may harm the training of the metanetwork. This step isn’t necesary, but facilitates convergence, saves VRAM use, and helps achieve better pruning results (Figure 8(b)) during most of the time.
(a) The origin network is easy to prune, after initial pruning and slight finetuning, it becomes hard to prune.
(b) Initial pruning helps get better metanetworks and better pruning results
Figure 8: With or without initial pruning (ResNet56 on CIFAR10 as an example)
Metanetwork: Change the origin hard to prune network into a graph, feed it through the metanetwork to get a new graph, and change it back into a new network.
Pruning: Prune the new easy to prune network.
Finetuning: Train the network with a relatively small learning rate. This is a common step that is widely used in pruning tasks.
Unlike prior learning to prune methods that require special training during each pruning, our pruning pipeline only requires a feed forward through metanetwork and standard finetuning, which is simple, general and effective.
Once we have a metanetwork, we can prune as many networks as we want. All our pruning models are completely different from meta-training data models. This demonstrates our metanetwork naturally has great transferability and no prior work has done something like this before.
Appendix D Experimental on CNNs
D.1 General Experiment Setup
D.1.1 General Settings
Data for Meta-Training: As discussed in Section 4.2, meta-training requires two types of data: data models and traditional training datasets. To generate a data model, we train a base model, perform initial pruning followed by finetuning, and finally save the model along with relevant information. Typically, generating 2–10 data models is sufficient, with 1–8 used for meta-training and the remainder reserved for meta-evaluation or testing.
Visualizing Metanetworks: This is a key technique used throughout our experiments. When we refer to visualizing a metanetwork, we mean passing a model through the metanetwork, followed by finetuning, and then visualize the “Acc VS. Speed-Up” curve of the resulting model, as illustrated in Figure 5. This visualization provides insights into the metanetwork’s behavior and helps determine whether it is suitable for pruning. An ideal metanetwork should exhibit a long flat region in the curve where accuracy remains close to or above the target pruning accuracy. There are two ways to select a suitable metanetwork: (1) using meta-evaluation during meta-training, or (2) visualizing metanetworks post-training. In our final experiments, we exclusively use the second method, which is significantly more efficient. We typically do not visualize all metanetworks but instead search for the best one using a binary search strategy.
D.1.2 General Meta Training Details
Equivalent Conversion between Network and Graph: During meta-training, we first convert the original network into an origin graph. Then, we feed this origin graph through the metanetwork to generate a new graph. Finally, we convert the new graph back into a new network. When converting the original network into the origin graph, all network parameters are mapped to either node or edge features in the graph. However, some graph features do not correspond to any network parameter. For these missing features, we use default values that effectively simulate the absence of those parameters. When converting the new graph back into the new network, we only map those trainable parameters from the new graph back to the new network. For untrainable parameters in the new network, we keep it the same as the old network. We provide more explanations for this conversion process in the following subsections: D.2, D.3, and D.4.
Meta Training Milestone: As discussed in the Ablation Study (Section 5.3), as the number of training epochs increases, the metanetwork’s ability to make the network easier to prune becomes stronger, while its ability to maintain the accuracy becomes weaker. During meta-training, we usually set a milestone for our learning rate scheduler. Before this milestone, we use a relatively large learning rate to quickly improve the metanetwork’s ability to make networks easier to prune. After the milestone, we switch to a smaller learning rate to finetune the metanetwork and slow down the change of the two abilities, allowing us to select a well-balanced metanetwork. To determine a reasonable milestone value, we can first perform a short meta-training phase using the large learning rate, then visualize the resulting metanetworks to identify the point where the accuracy-maintaining ability of metanetwork is a bit stronger than we expect.
D.1.3 General Pruning Details :